Three kids are tugging at a blanket with the following forces: 75N [E20ºN], 30N [W] and
100N [S30ºE]. Could you please help me find the magnitude of the net force on the blanket and direction in which its moving.

Thank you.
1
Expert's answer
2020-04-06T09:13:56-0400

The net force

"\bf F=F_1+F_2+F_3."

In projections on the axes

"F_x=F_1 \sin20\degree-F_2+F_3 \cos30\degree\\
F_y=F_1 \cos20\degree+0-F_3 \sin30\degree"

Hence

"F_x=75 \sin20\degree-30+100 \cos30\degree=82.3 N,\\
F_y=75 \cos20\degree-100 \sin30\degree=20.5 N."

The magnitude of a net force

"F=\sqrt{F_x^2+F_y^2}=\sqrt{82.3^2+20.5^2}=84.8 N."

The direction angle

"\theta =\tan^{-1}\frac{F_y}{F_x }=\tan^{-1}\frac{20.5}{82.3}=14\degree."

Answer: "84.8\rm N[N14\degree E]."